OpenStudy (anonymous):

True or False: If an exponential term of any base is raised to a power of 0, then the answer is zero. Briefly explain your answer.

OpenStudy (anonymous):

= 1 0 is false

OpenStudy (anonymous):

we can show this by an example x^6 / x^6 = 1 = x^(6-6) = x ^0 so x^0 = 1 and this is true for any base

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

yw